Given production function=(k+2)(l+1),cost of capital=$4,wage=$6.money available to the firm =8130, how to find q*?
1
Expert's answer
2021-10-15T11:23:06-0400

Production function : "q=(k+2)(l+1)"

"q=kl+k+2l+2"

Cost function: wage"\times" labor+cost of capital"\times" capital=Money available

"C:6l+4k=8130"

The firm would like to maximize the production that is the value of q with respect to its cost function:

We will use the lagrangian function:

Max q subject to C

"L=kl+k+2l+2-\u03bb(6l+4k-8130)"

Differentiating L wrt k, l, and λ

"\frac{dL}{dl}=k+2-6λ=0\\

k+2=6λ....(1)\\

\frac{dL}{dk}=l+1-4λ=0\\

l+1=4λ....(2)\\

\frac{dL}{dλ}=6l+4k-8130=0\\

6l+4k=8130....(3)\\"

 

"From (1) and (2)\\

\frac{(k+2)}{(l+1)}=\frac{3}{2}\\

\frac{2(k+2)}{3}=l+1\\

\frac{2k}{3}+\frac{4}{3}-1=l\\

\frac{2k}{3}+\frac{1}{3}=l\\

Putting\space this\space in \space (3)\\

6l+4k=8130\\

6(\frac{2k}{3}+\frac{1}{3})+4k=8130\\

4k+2+4k=8130\\

8k=8128\\

k=1016"

Putting the value of k in the relation to get l

"\frac{2k}{3}+\frac{1}{3}=l\\

2k+1=3l\\

2(1016)+1=3l\\

l=677.67"

To get the value of q, we will put the obtained optimal value of k and l in the production function:

"q=(k+2)(l+1)

q=(1016+2)(677.67+1)\\

q=1018\times 678.67\\

q=690,886.06"